15:15 Start Shift working with low current Moller Runs as well as other work on daq and event rate.
16:50 Noticed the BTU MCC statistics link is not functioning, so will have to make a estimate for MCC.
20:00 Sasha has completed Moller studies
20:10 Request controlled access to Hall to check the third arm threshold settings.
21:10 Request beam at 100nA with raster settings fast: 2X2 mm x-low:4.023 x-high:5.455 y-low:4.082 y-high:5.522
21:30 Requesting 280nA test of tungsten calorimeter (Dave M. calculation yields that the target setup can not withstand 100nA with LHe drained so target will remain full while Jixie decided next step.)
22:10 Fast raster off requesting 100nA
22:25 Requesting 1500nA pulsed for testing
22:35 Requesting 150nA cwb
23:10 target moved to carbon 12-40 with both rasters off at 100nA
23:16 Change beam position to 0, 0 for harp scan
23:36 Moved back to crosshair at 3micA pulsed